# Explaining mountain pine beetle dynamics: From life history traits to large, episodic outbreaks

### v1

@@ 5 Acknowledgements
We would like to thank Micah Brush and Antonia Musso for insightful discussions; and Xiaoqi Xie for obtaining the Biosim data (Figs. D.5 & D.4. This research was funded through a grant to the TRIA-FoR Project to ML from Genome Canada (Project No. 18202).
6 Data and code availability statement
All code and almost all data will be made available shortly, both on Github and a TBD data repository.
When seeking data from publications, we initially searched for downloadable data files. Where these weren’t available, we used the GraphGrabber software and Microsoft Excel’s digitizing features to extract data from figures and tables. Fixed-wing Aerial Detection Survey data for the western United States was downloaded from various (region-specific) USDA Forest Service websites. Fixed-wing Aerial Overview survey data for British Columbia was downloaded from the BC forestry FTP site (as of December 12, 2023, https://www.for.gov.bc.ca/ftp/). Heli-GPS Aerial Survey data for Alberta — here utilized in Figs. 1 & 2 — was provided by the provincial government of Alberta. However, this data cannot be made publicly available per our data-sharing agreement.
@@ Appendix A Statistical model description
A.1 Overview
@@
There are three statistical models. 1) The tree growth model (Appendix A.2); 2) The simple Allee effect model (Appendix A.6), which is fit to the Boone et al., dataset; and the complex outbreak model (Appendices A.3A.7), which is fit to the Klein et al., dataset. Table A.1 describes all relevant notation, including subscripts, data, intermediate quantities, and bonafide parameters.
In the complex outbreak model, each life-history process is described in its own sub-model (Appendices A.3A.6). Going further, the deterministic skeletons of the sub-models are strung together to form the integrator model (Appendix A.7). This model integrates over unobserved quantities (e.g., beetles remaining after dispersal) to estimate additional model parameters.
@@
In the main text, we presented the simulation model with each time-step beginning and ending with the density of emerging beetles. This approach was chosen for two reasons: Firstly, flying beetles are a noticeable aspect of actual outbreaks. Secondly, emergence ties directly into our discussion of the interaction between peak beetle density and long-distance dispersal.
In the outbreak statistical model, each time step begins and ends with beetles that have successfully mass attacked a tree. This model structure gives us an additional year of data at the beginning of the outbreak; the later years of the outbreak are dominated by stand-dependent dispersal, and are therefore less informative with respect to the estimation of many non-dispersal parameters.
A.2 Lodgeple pine forest dynamics
@@
Residual errors in the densities of small and large-diameter trees are made commensurate through the logarithmic transformation. The noise parameter is $\sigma_{\text{tree}}$. The model provides a good fit to the data (Fig. A.1), though there is substantial uncertainty due to the high quotient of parameters to data (3 and 7 respectively).
@@
A.3 Size-dependent preference sub-model
@@
High values of preference indicate that beetles like to attack a DBH class disproportionately, in relation to the surface area that is “owned” by that DBH class. The preference variable can be calculated directly from the Klein et al., data. The data show that beetle preference increases linearly with tree diameter, and that there is clear heteroskedasticity that is explained by small sample sizes for larger DBH classes (Fig. 10)
@@
where $\sigma_{\text{{pref0}}}$ accounts for a baseline level of variability, and $\sigma_{\text{{pref1}}}$ accounts for increasing variability with smaller sample sizes, with the typical inverse square root scaling that would be expected under the central limit theorem. We do not know the exact sample sizes, so we treat the total tree surface area (within the $j^{th}$ diameter class) as a proxy for sample size.
A.4 Converting beetle attacks to tree deaths
We model killed trees as a linear regression through the origin, with a dynamic noise term to account for some obvious heteroskedasticity (Fig. 12). The killed trees per acre, denoted $(KTPA)_{j,y}$ is normally distributed:
@@ A.4 Converting beetle attacks to tree deaths
A.5 Size-dependent fecundity
Figure A.2: Observations vs. predictions for the size-dependent fecundity sub-model. The error bars give the 95% predictive intervals, representing both parameter and sampling uncertainty. The dashed grey line diagonal line represents perfect agreement between observed and predicted values. Data from Klein et al., (1978).
The Klein et al., dataset does not contain DBH-specific emergence data. The total number of emergence holes per acre is normally distributed
@@ A.5 Size-dependent fecundity
A.6 Allee effect
The Boone et al., dataset contains two pertinent variables: density of trees attacked and proportion of attacked trees that are killed. As a first step, we convert the number of attacked trees to the number of attacking beetles. This is done primarily for practical reasons, as we aim to model the Klein et al., data, which lacks information on the number of attacked trees. Additionally, from a biological standpoint, using the number of attacked trees as a success indicator could be misleading. For instance, a scenario where numerous beetles attack a single large tree would erroneously imply a low success rate. This contradicts observations that beetles are more successful when they concentrate on larger trees.
To estimate the number of attacking beetles, we multiply the attacked trees by 360, the average number of attacks per successfully attacked tree in the Klein et al.,. Since successfully attacked trees receive more attacks, this is probably an overestimate of beetles per attacked tree, which would ultimately lead to a less-informative prior for the Allee threshold parameter, $a$, when modelling the Klein et al., data.
@@ A.6 Allee effect
Recall that the Allee effect sub-model is first fit to the Boone et al., data (Fig. 8). The subsequent estimate of $a$ is used to set an informative prior for the outbreak model (which is constituted by sub-models and the integrator model, which is then fit to the Klein et al., data. Specifically, we choose a half-normal prior with a standard deviation equal to half of the posterior mean of $a$.
The normal distribution is unrealistic because it implies that probabilities outside of the interval $[0,1]$ are possible. The normal distribution assumes constant variance, whereas the Boone et al., data clearly shows heteroskedasticity. That being said, a more realistic error structure would unnecessarily complicate the model. The normality of errors and constant variance are unimportant assumptions when trying to estimate the conditional expectation (Gelman and Hill,, 2007, Sec. 3.6), which is exactly what we are trying to do — only the mean probability of success (i.e., Eq. 32) is used to model the Klein et al., data.
A.7 Integrator model
The integrator model combines the deterministic skeletons of all sub-models (Appendices A.3A.6) to predict the next year’s gallery starts, solely from this year’s gallery starts. In doing so, we can estimate several parameters that cannot be estimated from singular sub-models: $a$, $s_{0}$, $\lambda_{0}$, $\lambda_{1}$, and a residual noise parameter, $\sigma_{\text{resid}}$.
@@ A.7 Integrator model
To predict the number of attacking beetles, we multiply the exit hole density by the predicted dispersal-phase survival. Contrary to Eq. 4 in the simulation model, there is never any immigration:
@@
To complete the virtual life-cycle and predict the number of successful attacks next year, denoted $\mu_{y,\text{full}}$, the Allee effect is applied to the predicted attack density
$\mu_{y+1,\text{full}}=\hat{(\text{{attacks per acre}})_{y+1}}\times\left(1-\exp\left(-2.3\cdot\frac{{\hat{(\text{{attacks per acre}})_{y+1}}}}{a}\right)\right)$ (38)
@@
$t$
Tree age in years
N/A
Lodgepole, A.2
j
The DBH-class index, uniquely identifies a tree diameter class
N/A
All
y
The year index,
N/A
All
tree
Pertaining to the lodgepole pine dyanamics sub-model
N/A
Lodgepole, A.2
pref
Pertaining to the size-dependent attack preference sub-model
N/A
Attack pref., A.3
dead
Pertaining to the tree to killed tree sub-model
N/A
Killed trees, A.4
exit
Pertaining to the size-dependent fecundity sub-model
N/A
Fecundity, A.5
allee
Pertaining to the Allee effect sub-model
N/A
Allee effect, A.6
@@
Upper age limit of trees within a particular diameter class
years
Lodgepole, A.2
@@
Lower age limit of trees within a particular diameter class
years
Lodgepole, A.2
$\text{TPA}_{j}$
Density of trees within the ages corresponding to the $j^{th}$ diameter class
@@
Lodgepole, A.2
$(\text{{preference}})_{j,y}$
Mass attack preference, defined as the attack density within a particular diameter class, divided by the attack density across all trees (within a year)
unitless
Attack pref., A.3
$\text{{DBH}}_{j}$
Diameter at breast height
inches
Attack pref., A.3
$(\text{{KTPA}})_{j,y}$
Killed trees per acre
@@
Gallery starts per acre
@@
Killed trees, A.4
$(\text{surface area per tree})_{j}$
Susceptible surface area per individual tree, defined by Eq. 1
$\text{meters}^{2}\cdot$ $\text{trees}^{-1}$
Killed trees, A.4
$(\text{{exit density}})_{y}$
Emergence holes per acre
@@
Fecundity, A.5
$J$
Total number of unique diameter classes,
unitless
Fecundity, A.5
$Y$
Total number of years in the data set
unitless
Fecundity, A.5
$(\text{{attacks per acre}})_{j,y}$
Attacking female beetles per acre
@@
Proportion of total susceptible tree surface area belonging to a particular diameter class
unitless
Integrator, A.7
Derived quantities
$\mu_{j,\text{tree}}$
Expected tree density in the $j^{th}$diameter class
@@
Lodgepole, A.2
$\mu_{j,y,\text{{pref}}}$
Predicted mean of the mass attack preference
unitless
Attack pref., A.3
$\sigma_{j,y,\text{{pref}}}$
Standard deviation of the mass attack preference
unitless
Attack pref., A.3
$\mu_{j,y,\text{{dead}}}$
Predicted mean of the killed trees density
@@
Killed trees, A.4
$\sigma_{j,y,\text{{dead}}}$
Standard deviation of the killed trees density
@@
Killed trees, A.4
$\mu_{y,\text{{exit}}}$
Predicted mean of emergence hole density
(exit $(\text{exit holes})\text{acres}^{-1}$
Fecundity, A.5
$\mu_{j,y,\text{{allee}}}$
Mean proportion of successful attacks
unitless
Allee effect, A.6
$s_{y}$
Actual survival probability for beetles during the dispersal phase
unitless
Integrator, A.7
$\lambda_{y}$
Stand-dependent dispersal factor; actual mortality increases as the proportion of remaining susceptible surface area decreases
unitless
Integrator, A.7
$\mu_{y,\text{full}}$
Expected density of successful beetle attacks (across all DBH classes)
@@
Integrator, A.7
Model parameters
$K$
Tree carrying capacity
@@
Lodgepole, A.2
$m$
Annual probability of tree death (not caused by MPB)
unitless
Lodgepole, A.2
$\sigma_{\text{tree}}$
Noise parameter
@@
Lodgepole, A.2
$\beta_{0}$
Y-intercept of the piecewise linear preference-DBH function
unitless
Attack pref., A.3
$\beta_{1}$
Slope of the piecewise linear preference-DBH function; the DBH threshold for attacks is $-\beta_{0}/\beta_{1}$
unitless
Attack pref., A.3
$\sigma_{\text{{pref0}}}$
noise parameter; accounts for a baseline level of variability in mass attack preference
unitless
Attack pref., A.3
$\sigma_{\text{{pref1}}}$
Noise parameter; accounts for increasing variability with smaller sample sizes in mass attack preference
unitless
Attack pref., A.3
$\gamma$
Proportionality constant for number of killed trees; the reciprocal of the average attack density
$\text{meters}^{2}\cdot$ $(\text{gallery starts})^{-1}$
Killed trees, A.4
$\sigma_{\text{{dead0}}}$
noise parameter; the effect of the relative attack density on the standard deviation
$\text{meters}^{2}\cdot$ $(\text{gallery starts})^{-1}$
Killed trees, A.4
$\zeta_{0}$
Y-intercept of the piecewise linear emergence-DBH function
@@
Fecundity, A.5
$\zeta_{1}$
Slope of the diameter-emergence regression; the DBH threshold for emergence is $-\zeta_{0}/\zeta_{1}$
@@
Fecundity, A.5
$\sigma_{\text{{exit}}}$
Standard deviation parameter for emergence holes
@@
Fecundity, A.5
$a$
Allee threshold (density of beetles at which probability of attack success is 0.9)
@@
Allee effect, A.6
$\sigma_{\text{{allee}}}$
Standard deviation parameter for proportion of successful attacks
unitless
Allee effect, A.6
$s_{0}$
Maximum survival probability for beetles during the dispersal phase
unitless
Integrator, A.7
$\lambda_{1}$
Sensitivity of dispersal mortality on tree surface area; controls the speed of the transition from low to high dispersal mortality as trees are killed.
unitless
Integrator, A.7
$\lambda_{0}$
Transition point parameter; controls the level of tree death at which dispersal mortality becomes significant
unitless
Integrator, A.7
$\sigma_{\text{resid}}$
Residual noise parameter, to account for error in the predicted density of successful beetle attacks next year

### v2

@@ 5 Acknowledgements
We would like to thank Micah Brush for insightful discussions; and Xiaoqi Xie for obtaining the Biosim data (Figs. B.7 & B.6). This research was funded through a grant to the TRIA-FoR Project to ML from Genome Canada (Project No. 18202).
@@ Appendix A Statistical model description
A.1 Data
Klein et al.,’s study was conducted from 1965–1972 in the southwest corner of Yellowstone National Park in Idaho. Data was collected and aggregated across 85 permanent plots that spanned a square mile. Trees were grouped into six 2-inch diameter classes, ranging from 8 to $\geq$ 16 inches. For each of the 106 sampled trees, paired 6x6-inch sample squares were marked on opposite sides of the bole at regular height intervals. Within these squares, emergence holes and successful attacks (operationalized as egg-gallery starts) were measured (Fig. A.1). Some of this data was originally presented by Parker, (1973), but a more thorough reporting is given by Klein et al., (1978).
Figure A.1: Time series of emergence holes and gallery starts in the Klein et al., dataset. Emergence and subsequent attacks are assumed to take place in July and August respectively. There are always far fewer gallery starts than emergence holes, suggesting high levels of dispersal mortality.
A.2 Model overview
@@
There are three statistical models. 1) The tree growth model (Appendix A.4); 2) The simple Allee effect model (Appendix A.8), which is fit to the Boone et al., dataset; and 3) the complex outbreak model (Appendices A.5A.9), which is fit to the Klein et al., dataset. Table LABEL:tab:model_notation describes all relevant notation, including subscripts, data, intermediate quantities, and bonafide parameters.
The tree growth model (Appendix A.4) estimates the demographic parameters of lodgepole pine: density-independent mortality and forest carrying capacity ($m$ and $K$). This model was fit with the 1965 pre-outbreak tree diameter distribution in the Klein et al., dataset (sample size = 7). The simple Allee effect model (Appendix A.8) describes the Allee effect and was fit with Boone et al., data (sample size = 28). The posterior distribution of the Allee threshold parameter was used to create an informative prior for the Allee effect in our third statistical model.
The third statistical model (Appendices A.5A.9), which is by far the most complex, uses the Klein et al., dataset (sample size = 91) to model the outbreak dynamics of MPB. This outbreak model contains several sub-models, simple models that are used to resolve one aspect of MPB’s life cycle. For example, tree diameter and beetle attack preference, are used to estimate the parameters $\beta_{0}$ and $\beta_{1}$. A variety of such sub-models provide direct estimates of the parameters $\beta_{0}$, $\beta_{1}$, $\gamma$, $\zeta_{0}$, $\zeta_{1}$, and additional noise parameters. The remaining parameters ($a$, $s_{0}$, $\lambda_{0}$, and $\lambda_{1}$), are estimated by stringing together the deterministic aspects of the sub-models to predict the number of gallery starts in the following year; we call this the integrator model. All of the sub-models, along with the integrator model are combined and fit simultaneously. The result is a hierarchical model structure, where information is partially pooled across sub-models, because all parameters must work cohesively to predict next year’s gallery starts.
@@
In the main text, we presented the simulation model with each time-step beginning and ending with the density of emerging beetles. This approach was chosen for two reasons: Firstly, flying beetles are a noticeable aspect of actual outbreaks. Secondly, emergence ties directly into our discussion of the interaction between peak beetle density and long-distance dispersal. In contrast, the outbreak statistical model ends each step with beetles that have successfully mass-attacked a tree. This structure yields an extra year of data at the beginning of the outbreak’s. Later years, dominated by stand-dependent dispersal, offer less insight into non-dispersal parameters.
Throughout this appendix, we will refer to the surface area over which beetles may attack as the susceptible surface area; note that beetles do not attack the upper part of the bole due to thin phloem. To convert between tree diameter and susceptible surface area, we use a simple linear model,
$SA=-3.792+0.821\times DBH,$ (16)
A.3 Model-fitting
The parameters of the simulation model are estimated with three statistical models: the lodgepole-dynamics model, the Allee-effect model, and the outbreak-dynamics model. The combined statistical models are conceptually identical to the simulation model.
The third statistical model (Appendices A.5A.9), which is by far the most complex, uses the Klein et al., dataset (sample size = 91) to model the outbreak dynamics of MPB. This outbreak model contains several sub-models, simple models that are used to resolve one aspect of MPB’s life cycle. For example, tree diameter and beetle attack preference, are used to estimate the parameters $\beta_{0}$ and $\beta_{1}$. A variety of such sub-models provide direct estimates of the parameters $\beta_{0}$, $\beta_{1}$, $\gamma$, $\zeta_{0}$, $\zeta_{1}$, and additional noise parameters. The remaining parameters ($a$, $s_{0}$, $\lambda_{0}$, and $\lambda_{1}$), are estimated by stringing together the deterministic aspects of the sub-models to predict the number of gallery starts in the following year; we call this the integrator model. All of the sub-models, along with the integrator model are combined and fit simultaneously. The result is a hierarchical model structure, where information is partially pooled across sub-models, because all parameters must work cohesively to predict next year’s gallery starts.
A.4 Lodgeple pine forest dynamics
@@
Residual errors in the densities of small and large-diameter trees are made commensurate through the logarithmic transformation. The noise parameter is $\sigma_{\text{tree}}$. The model provides a good fit to the data (Fig. A.2), though there is substantial uncertainty (Fig. A.3) due to the high quotient of parameters and data (i.e., $3/7$).
Figure A.2: The pre-outbreak diameter distribution of lodgepole pine is effectively modeled by assuming density-independent mortality and instant recruitment. The theoretical distribution of trees per acre in the $j^{th}$ diameter class is $(\text{TPA})_{j,\infty}=(1-m)^{(j-1)}mK$, so the logarithm of trees per acre within should be approximately linear with respect to diameter at breast height (DBH). Data from Klein et al., (1978). The predictions were generated with posterior means of model parameters.
@@
A.5 Tree size-dependent preference sub-model
@@
High values of preference indicate that beetles like to attack a DBH class disproportionately, in relation to the surface area that is “owned” by that DBH class. The preference variable can be calculated directly from the Klein et al., data. The data clearly show that that preference increases linearly with tree diameter (Fig. A.4, A.5).
Figure A.4: Beetles disproportionately kill large trees. The x-axis is the proportion of susceptible surface area belonging to live trees of a particular diameter class, within a year. The y-axis is the proportion of total attacks received by the same diameter class, within a year. The black line is a LOESS model prediction. The grey dashed line is the one-to-one line, representing the no-mass-attack-preference hypothesis. Data from Klein et al., (1978).
Figure A.5: Beetle preference is an increasing function of tree diameter. The data (from Klein et al.,, 1978) is well-modeled by a piecewise linear function (see Eq. 22). There is clear heteroskedasticity in the data, which is likely due to smaller sample sizes within classes of large-diameter trees. The exact number of samples is not available, so we use the within-DBH-class surface area as a proxy for sample size, visualized as scatter-plot point size. The prediction was generated using posterior means of model parameters.
@@
where $\sigma_{\text{{pref0}}}$ accounts for a baseline level of variability, and $\sigma_{\text{{pref1}}}$ accounts for increasing variability with smaller sample sizes (clearly visible in Fig. A.5), with the typical inverse square root scaling that would be expected under the central limit theorem. We do not know the exact sample sizes, so we treat the total tree surface area (within the $j^{th}$ diameter class) as a proxy for sample size.
A.6 Converting beetle attacks to tree deaths
We model killed trees as a linear regression through the origin, with a dynamic noise term to account for some obvious heteroskedasticity (Fig. A.6). The killed trees per acre, denoted $(KTPA)_{j,y}$ is normally distributed:
@@ A.6 Converting beetle attacks to tree deaths
Figure A.6: The number of killed trees is a linear function of successful beetle attacks (corrected for DBH-specific surface area), suggesting that beetles “fill up” a tree to a universal density, and then move on to the next tree. The prediction was generated using posterior means of model parameters. Data from Klein et al., (1978).
A.7 Tree size-dependent fecundity
The Klein et al., dataset does not contain DBH-specific emergence data. The total number of emergence holes per acre is normally distributed:
@@ A.7 Tree size-dependent fecundity
Figure A.7: Observations vs. predictions for the tree size-dependent fecundity sub-model. The error bars give the 95% predictive intervals, representing both parameter and sampling uncertainty. The dashed grey line diagonal line represents perfect agreement between observed and predicted values. Data from Klein et al., (1978).
We cannot justify the functional form of Eq. 28 with graphical evidence, since the Klein et al., dataset only contains total emergence holes (not stratified by diameter class). However, the linear nature of the relationship has been well-established (Reid,, 1963, Table 1; Safranyik,, 1968, Fig. 25; Cole,, 1969, Fig. 5; Cudmore et al.,, 2009, Fig. 6). The linear relationship also implicitly accounts for the fact that the male:female sex ratio of emerging beetles increases with tree diameter (Cole et al.,, 1976, Fig. 5). Note that emergence only depends on killed trees, not attacking beetles, because a slightly above-optimal attack density is balanced by intraspecific density dependence, manifesting as a decrease in per capita oviposition or early-instar survival. In other words, there is perfectly compensatory density dependence at the tree scale (Fig. E.3; Berryman,, 1974, Fig. 2c; Berryman,, 1976, Fig. 2).
A.8 Allee effect
The Boone et al., dataset contains two pertinent variables: density of trees attacked and proportion of attacked trees that are killed. As a first step, we convert the number of attacked trees to the number of attacking beetles. This is done primarily for practical reasons, as we aim to model the Klein et al., data, which lacks information on the number of attacked trees (only successfully attacked trees are reported)
To estimate the number of attacking beetles, we multiply the attacked trees by 360, the average number of attacks per successfully attacked tree in the Klein et al.,. Since successfully attacked trees receive more attacks, this is probably an overestimate of beetles per attacked tree, which would ultimately lead to a less informative prior for the Allee threshold parameter, $a$, when modeling the Klein et al., data.
@@ A.8 Allee effect
The exponential saturating function above is justified by the Boone et al., dataset (Fig. A.8). The Allee effect sub-model is first fit to the Boone et al., data; the subsequent estimate of $a$ is used to set an informative prior for the full statistical outbreak model (constituted by the Allee effect sub-model, other sub-models, and the integrator model), which is fit to the Klein et al., data. Specifically, we choose a half-normal prior with a standard deviation equal to half of the posterior mean of $a$.
Figure A.8: The probability of an attacked tree being killed is an increasing function of beetle density. The data is effectively modeled by an exponential saturating function (see Eq. 3, “Probability of success” term). The prediction was generated using posterior means of model parameters. Data from the Boone et al., dataset.
The normal distribution in Eq. 29 is unrealistic because it implies that probabilities outside of the interval $[0,1]$ are possible. Additionally, the normal distribution assumes constant variance, whereas the Boone et al., data clearly shows heteroskedasticity. That being said, a more realistic error structure would unnecessarily complicate the model. The normality of errors and constant variance are the least important assumptions when trying to estimate the conditional expectation (Gelman and Hill,, 2007, Sec. 3.6). This is exactly what we are trying to do: only the mean probability of success (i.e., Eq. 30) is used to model the Klein et al., data.
Another unrealistic aspect of the Allee effect sub-model is that the probability of success approaches an asymptote of 100%, whereas visual inspection of the data suggests an approximate maximum of 80-90% Fig. A.8. However, in the integrator model (Section A.9), an additional parameter for the maximum probability of successful attacks would be non-separable with the maximum dispersal-phase survival, $s_{0}$. Thus, our estimate of $s_{0}$ not only represents maximum survival during the dispersal phase, but also the fact that some mass attacks are unsuccessful despite high beetle density. Additionally, $s_{0}$ captures the phenemenon where low-vigor females successfully attack trees but do not oviposit (5–30% of females could fall into this category; Amman,, 1975, 1980).
A.9 Integrator model
The integrator model combines the deterministic skeletons of all sub-models (Appendices A.5A.8) to predict the next year’s gallery starts, solely from this year’s gallery starts. In doing so, we can estimate several parameters that cannot be estimated from singular sub-models: $a$, $s_{0}$, $\lambda_{0}$, $\lambda_{1}$, and a residual noise parameter, $\sigma_{\text{resid}}$.
@@ A.9 Integrator model
The equation for successful attacks, Eq. 4, shows how beetles are distributed among tree diameter classes, but not how beetles are distributed among individual trees. As it turns out, killed trees receive an approximately invariant number of attacks per unit surface area (Fig. A.9; see further discussion in Section 1.2.3).
Figure A.9: Successful attack density (gallery starts per square meter) increases slightly over time, but does not depend on tree diameter at breast height (DBH). Data from Klein et al., (1978).
If beetles are unsuccessful, the tree survives and the beetles are assumed to die. This is a simplifying assumption because unsuccessful beetles could attack other trees. The assumption is perhaps justifiable because some beetles will be killed by the first tree, the remainder face increased mortality risk during secondary dispersal, and late-arriving beetles have low fitness (Raffa,, 2001).
@@
To predict the number of attacking beetles, we multiply the exit hole density by the predicted probability of survival during the dispersal phase. This quantity is a logistically increasing function (from 0 to the maximum $s_{0}$) of “relative surface area” which is defined as the susceptible surface area of all trees with $DBH\geq 8\,\text{inches}$ in year $y$, divided by the total susceptible surface area when MPB is absent and the forest is at equilibrium. We use the relative surface area mostly for interpretability (e.g., 0.5 means the forest is half-depleted) but also for computational reasons (the Bayesian model-fitting program Stan prefers all parameters to be on the unit-scale). Unlike the simulations, there are no stochastic influxes of beetles:
@@
We cannot directly justify the functional form of Eq. 35 — say, by plotting total emergence against subsequent total attacks — because Klein et al., only reported attacks on successfully killed trees. However, we note that the logistic function is a standard way to model probabilities, and that the extreme limit of the function (i.e., dispersal mortality is near 100% when there are few trees) is supported by both the Klein et al., dataset (Fig. E.4) and the literature (see Section 1.2.2). In particular, Powell and Bentz, (2014) estimated that mean dispersal distance increased from several meters to several kilometers as tree density decreased.
To complete the virtual life-cycle and predict the number of successful attacks next year, denoted $\mu_{y,\text{full}}$, the Allee effect is applied to the predicted attack density:
$\mu_{y+1,\text{full}}=\hat{(\text{{attacks per acre}})_{y+1}}\times\left(1-\exp\left(-2.3\cdot\frac{{\hat{(\text{{attacks per acre}})_{y+1}}}}{a}\right)\right).$ (36)
@@
$t$ Tree age in years N/A Lodgepole, A.4
j The DBH-class index, uniquely identifies a tree diameter class N/A All
y The year index, N/A All
tree Pertaining to the lodgepole pine dyanamics sub-model N/A Lodgepole, A.4
pref Pertaining to the tree size-dependent preference sub-model N/A Attack pref., A.5
dead Pertaining to the tree to killed tree sub-model N/A Killed trees, A.6
exit Pertaining to the tree size-dependent fecundity sub-model N/A Fecundity, A.7
allee Pertaining to the Allee effect sub-model N/A Allee effect, A.8
@@
$(\text{{preference}})_{j,y}$ Mass attack preference, defined as the attack density within a particular diameter class, divided by the attack density across all trees (within a year) unitless Attack pref., A.5
$\text{{DBH}}_{j}$ Diameter at breast height inches Attack pref., A.5
@@
$(\text{surface area per tree})_{j}$ Susceptible surface area per individual tree, defined by Eq. 16 $\text{meters}^{2}\cdot$ $\text{trees}^{-1}$ Killed trees, A.6
@@
$J$ Total number of unique diameter classes, unitless Fecundity, A.7
$Y$ Total number of years in the data set unitless Fecundity, A.7
@@
$\mu_{j,y,\text{{pref}}}$ Predicted mean of the mass attack preference unitless Attack pref., A.5
$\sigma_{j,y,\text{{pref}}}$ Standard deviation of the mass attack preference unitless Attack pref., A.5
@@
$\mu_{y,\text{{exit}}}$ Predicted mean of emergence hole density (exit $(\text{exit holes})\text{acres}^{-1}$ Fecundity, A.7
$\mu_{j,y,\text{{allee}}}$ Mean proportion of successful attacks unitless Allee effect, A.8
$s_{y}$ Actual survival probability for beetles during the dispersal phase unitless Integrator, A.9
@@
Model parameters
@@
$m$ Annual probability of tree death (not caused by MPB) unitless Lodgepole, A.4
@@
$\beta_{0}$ Y-intercept of the piecewise linear preference-DBH function unitless Attack pref., A.5
$\beta_{1}$ Slope of the piecewise linear preference-DBH function; the DBH threshold for attacks is $-\beta_{0}/\beta_{1}$ unitless Attack pref., A.5
$\sigma_{\text{{pref0}}}$ noise parameter; accounts for a baseline level of variability in mass attack preference unitless Attack pref., A.5
$\sigma_{\text{{pref1}}}$ Noise parameter; accounts for increasing variability with smaller sample sizes in mass attack preference unitless Attack pref., A.5
$\gamma$ Proportionality constant for number of killed trees; the reciprocal of the average attack density $\text{meters}^{2}\cdot$ $(\text{gallery starts})^{-1}$ Killed trees, A.6
$\sigma_{\text{{dead0}}}$ noise parameter; the effect of the relative attack density on the standard deviation $\text{meters}^{2}\cdot$ $(\text{gallery starts})^{-1}$ Killed trees, A.6
@@
$\sigma_{\text{{allee}}}$ Standard deviation parameter for proportion of successful attacks unitless Allee effect, A.8
$s_{0}$ Maximum survival probability for beetles during the dispersal phase unitless Integrator, A.9
$\lambda_{1}$ Sensitivity of dispersal mortality on tree surface area; controls the speed of the transition from low to high dispersal mortality as trees are killed. unitless Integrator, A.9
$\lambda_{0}$ Transition point parameter; controls the level of tree death at which dispersal mortality becomes significant unitless Integrator, A.9
@@
Appendix B Additional model justification
B.1 Stochastic perturbation size
To simulate the influx of female beetles corresponding to the incipient-epidemic transition, we set $I=300$ (see Eq. 2 in the main text). This number was obtained with a back-of-the-envelope calculation: the median emergence density of females across several studies and Forest Service reports (Other Studies in Fig. E.6), multiplied by the susceptible surface area of typical killed tree (10 inches DBH as an input to Eq. 16), multiplied by the median estimate of the dispersal-phase survival probability (using gallery-starts/emergence-holes quotient from the Klein et al., dataset, Fig. E.4), multiplied by 0.81, which is the number of successfully attacked trees per acre that corresponds to an incipient-epidemic state (according to Carroll et al., 2006c, ). The chosen values of $P_{I}$ and $I$ are educated guesses, but we re-ran our analysis with several different values and obtained qualitatively identical results.
B.2 Mass attack preference does not change
One potential shortcoming of our model is that preference only depends on tree diameter, whereas previous papers have shown that MPB’s preference for large-diameter trees increases as outbreaks progress (Boone et al.,, 2011; Howe et al., 2022b, ). To investigate further, we separately fit our preference sub-model to each year of data and plotted $\beta_{1}$ (the slope of the preference $\sim$ DBH regression) against beetle pressure, as measured by the density of emerging females, and the number of emerging females per unit of susceptible tree surface area (Fig. B.1). There was no clear pattern, indicating that dynamic changes to the DBH preference are small or nonexistent. Similarly, there was no relationship between beetle pressure and the residuals of our original preference model (Fig. B.2).